250. 0 mL of 1.00 M NaOH solution has a density of 1.90 g/mL. Calculate the concentration of this solution in molality.



1
Expert's answer
2022-01-04T01:25:03-0500

Molar mass of NaOH= 39.997

250/39.997=6.25

1/6.25=0.16

= 0.16/0.0019

= 84.21mol/Kg